OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f of x as x approaches negative 1 where f of x equals 4 minus x when x is less than negative 1, 5 when x equals negative 1, and x plus 6 when x is greater than negative 1

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

lol can you format it?

OpenStudy (phi):

can you use the equation editor to rewrite that?

OpenStudy (anonymous):

There is the question if that's ok

OpenStudy (anonymous):

@phi where you at tho

OpenStudy (phi):

to have a limit you want to get the same number as you approach x=-1 from below or from above. when x < -1 we use the first expression 4- x as x gets closer to -1 we get close to 4 -(-1) = 4+1 =5

OpenStudy (anonymous):

so the limit is 5?

OpenStudy (phi):

you also have to get 5 when x > -1 and you get closer to -1 so use the expression x+6 when x>-1 as x gets close to -1, what does x+6 get close to?

OpenStudy (anonymous):

5

OpenStudy (phi):

we get the same number from "above" and from "below" so we have a limit it is 5 (notice the middle expression 5 for x=-1 does not really matter)

OpenStudy (anonymous):

cool can you help with one more?